OpenStudy (anonymous):

Solve the linear system using subsitiution. Please show steps x = y+3 2x - y = 5 ?

OpenStudy (anonymous):

x = y+3 2x - y = 5 2(y+3) - y = 5 2y + 6 - y = 5 y = 5 - 6 so, y is equal to...?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

Great

hero (hero):

x = y + 3 ---> y = x - 3 2x - y = 5 --> y = 2x - 5 y = y x - 3 = 2x - 5 2x - x = 5 - 3 x = 2 y = -1

hero (hero):

Therefore, (x,y) = (2,-1)

OpenStudy (anonymous):

Thanks Hero .

hero (hero):

yw

hero (hero):

You finally made it to systems of equations. Only gets better from here

OpenStudy (anonymous):

Better as in harder or better as in easier ? 0.o

hero (hero):

Well, that depends on you. Everyone has a different take on it.

OpenStudy (anonymous):

Oh mmkay . We'll see .

OpenStudy (anonymous):

If you really want to learn, you will get it, isai. ^^